OpenStudy (anonymous):

Which expression is a factor of 3xy + 2x – 18y – 12? Answerss 2 + 3y 3x – 2y x + 6 x + 3

OpenStudy (king):

2+3y

OpenStudy (king):

3xy+2x-18y-12 =x(2+3y)-6(2+3y) =(x-6)(2+3y)

OpenStudy (king):

understood?